Sasha owns two investments, A and B, that have a combined total value of 46,900 dollars. Investment A is expected to pay 25,600 dollars in 3 year(s) from today and has an expected return of 12.63 percent per year. Investment B is expected to pay 45,107 dollars in T years from today and has an expected return of 3.68 percent per year. What is T, the number of years from today that investment B is expected to pay 45,107 dollars? Round your answer to 2 decimal places (for example, 2.89, 14.70, or 6.00).

1 year(s) ago, Mack invested 5,130 dollars. In 2 year(s) from today, he expects to have 7,810 dollars. If Mack expects to earn the same annual return after 2 year(s) from today as the annual rate implied from the past and expected values given in the problem, then how much does Mack expect to have in 5 years from today?
Answer #1

A). Combined investment = $46900

Investment A is expected to pay 25,600 dollars in 3 year(s) from today and has an expected return of 12.63 percent per year. Investment B is expected to pay 45,107 dollars in T years from today and has an expected return of 3.68 percent per year.

46900 = PV of investment A + PV of investment B

Using formula PV = FV/(1+r)^t

So, 46900 = 25600/(1+0.1263)^3 + 45107/(1+0.0368)^T

=> T = 12.24 years

B). 1 year(s) ago, Mack invested 5,130 dollars. In 2 year(s) from today, he expects to have 7,810 dollars.

So, PV = $5130

FV = $7810

t = 3 years

calculating r using PV=FV/(1+r)^t

so, 5130 = 7810/((1+r)^3

=> r = 15.04%

For 5 year from today it is 3 years from 2 year from today, So we can use

PV= 7810, r=15.04%, t=3years

So, FV = 7810*(1+0.1504)^3 = $11890

Mack expect to have $11890 in 5 years from today.